# On robustness of Spectral Rényi divergence

### v3

@@ 4 Simulation studies
    using the smoothed periodogram ($I^{S}_{n}$);
  • the minimum Itakura–Saito divergence estimator employing the periodogram ($I_{n}$); and
  • the minimum Itakura–Saito divergence estimator with the smoothed periodogram ($I_{n}^{S}$).
For each divergence minimization problem, the minimizer is computed via the gradient descent algorithm with a fixed learning rate $\gamma=0.005$ starting from a specified initial point detailed for each model. We stop the optimization when the optimization step reaches 10000 or when the Euclidean norm of the gradient of each divergence evaluated at each updated estimate is less than $10^{-3}$. For smoothing the periodogram, we apply the modified Daniell smoothers twice, where the length of the first time is $3$ and the length of the second time is $5$.
@@ 4.1 The Brune spectral model with attenuation
Figure 2: Spectral densities with the estimates plugged-in without any trend. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Figure 3: Spectral densities with the estimates plugged-in with the trigonometric trends. The red dashed lines denote the frequencies at which the outliers are injected. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Table 1: Mean values of biases with standard deviations without any trend. Values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
@@
IS with $I_{n}^{S}$ $\theta^{(0),2}$0.29 ($\pm$0.30) 1.95 ($\pm$1.89) 1.86 ($\pm$0.98)
R ($\alpha=0.50$) $\theta^{(0),3}$0.10 ($\pm$0.04) 0.51 ($\pm$0.13) -0.30 ($\pm$0.06)
R ($\alpha=0.75$) $\theta^{(0),3}$0.05 ($\pm$0.06) 0.08 ($\pm$0.19) 0.05 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),3}$0.007 ($\pm$0.06) -0.15 ($\pm$0.13) 0.72 ($\pm$0.35)
@@
IS with $I_{n}^{S}$ $\theta^{(0),3}$0.29 ($\pm$0.15) 2.09 ($\pm$1.68) 1.50 ($\pm$0.72)
Table 2: Mean values of biases with standard deviations with the trigonometric trends. Values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
@@
Consider the Brune spectral model [4] with attenuation [14] for velocity waveforms:
The Brune spectral model $\sigma^{2}/\{1+(\omega/\omega_{c})^{2}\}^{2}$ for displacement waveforms is the spectral density model of Matérn process [21] with the spectral decay rate fixed to 2 and often used in the earthquake source estimation [5, 33]. The attenuation factor $\exp(-\omega/Q)$ represents anelasticity of the medium through which the seismic wave propagates. The multiplicative factor $\omega^{2}$ comes from the time derivative that converts a displacement waveform to a velocity waveform.
@@
with $z_{1}=z_{2}=100$, $A_{1}=\pi/4$ and $A_{2}=\pi/8$. We report the results based on different values of $z_{1}=z_{2}$ in Appendix E. We use three different initial values: $\theta^{(0),1}=(1,0.1,1)$, $\theta^{(0),2}=(1,1,1)$, and $\theta^{(0),3}=(1,2,1)$. Setting $\sigma^{*}$ and $Q^{*}$ as initial values does not impact on the behavior of the minimum spectral Rényi divergence estimates; we report the results based on different initial values $\theta^{(0),4}=(0.1,0.1,0.1)$ and $\theta^{(0),5}=(2,2,2)$ in Appendix E.
Figures 2 and 3 display the estimated spectral densities. In each figure, the following representations are used:
@@
  • the solid blue curve illustrates the spectral density with the minimum Itakura–Saito divergence estimate plugged-in;
  • the solid red, salmon pink, and green curves display the spectral densities with the minimum spectral Rényi divergence ($\alpha=0.9,0.75,0.5$) estimates plugged-in, respectively.
@@
Tables 1 and 2 summaries the estimation results. Comparing the results based on different initial values, we find that the minimum Itakura–Saito divergence estimates are sensitive to the initial value of the optimization regardless of the existence of the smoother, while the spectral Rényi divergence yields estimation results stable to the choice of the initial value. The minimum Itakura–Saito divergence estimate with the smoothed periodogram performs quite worser than that with the periodogram. In the presence of outliers in the frequency domain, the minimum Rényi divergence estimate with $\alpha=0.5$ performs the best regardless of the choice of the initial value in this example.
@@ 5 Discussions
The value of $\alpha$ controls the trade-off between the the efficiency without the presence of outliers and the robustness even for larger trends. For an objective selection of $\alpha$, constructing an information criterion is one possible direction, but requires sufficient discussions and experiments. In practice, monitoring the behaviours of the minimum spectral Rényi divergence estimates for several values of $\alpha$ (say, $\alpha=0.5,0.75,0.9$) would be a workaround.
The minimum spectral Rényi divergence estimates can accommodate a superposition of outliers in the frequency domain, which is omitted in this paper because the extension is straightforward. This is also confirmed by numerical experiments in Section 4.
Numerical experiments presented in Section 4 suggest that the minimum spectral Rényi divergence estimates are robust to the choice of an initial value, while the minimum Itakura–Saito divergence estimate is sensitive to it. These robustness and sensitivity are confirmed in the other examples such as autoregressive models. The structure of the gradient $\mathcal{G}_{\alpha}(\theta\,;\,\widetilde{I}_{n})$ in the comparison with that of $\mathcal{G}_{1}(\theta\,;\,\widetilde{I}_{n})$ can give a clue to this, and so detailed analysis of these gradient landscapes would be one of interesting research directions.
6 Acknowledgement
The authors would like to thank Akifumi Okuno, Mirai Tanaka, Kei Kobayashi, Yuta Koike, Tomoyuki Higuchi, Hironori Fujisawa, Junho Yang, Masayuki Kano, and Shotaro Yagishita for their comments. This work is supported by JSPS KAKENHI (19K20222, 21H05205, 21K12067, 23K11024), MEXT (JPJ010217), and “Strategic Research Projects” grant (2022-SRP-13) from ROIS (Research Organization of Information and Systems). The simulation code is available at https://github.com/t-tetsuya/Spectral_Renyi.git.
@@ 6 Acknowledgement
  • Proof of Proposition 3 in Appendix D; and
  • Additional simulation studies in Appendix E.
@@ Appendix A Proof of Theorem 1
We firstly introduce the notation used in the proof. For each spectral density function $S$ of a stationary time series $\{X_{t}\}_{t\in\mathbb{Z}}$, we write $\Sigma_{n}(S)$ as the $n\times n$-Toeplitz matrix whose $(s,t)$th element is given by
@@
which is corresponding to the covariance matrix of $(X_{1},X_{2},\cdots,X_{n})^{\top}$. Then, the probability density function of $(X_{1},\cdots,X_{n})^{\top}$ is expressed by
@@
Let $\gamma\in(0,\infty)$. First introduce the $\gamma$-cross entropy $d_{\gamma}(p_{S},p_{\widetilde{S}})$ between two probability density functions $p_{S}$ and $p_{\widetilde{S}}$, which are not necessarily Gaussian probability density functions, is defined by
@@
Then the $\gamma$-divergence can be written as $D_{\gamma}[p_{S}:p_{\widetilde{S}}]=d_{\gamma}[p_{S}~{}:p_{\widetilde{S}}]-d_{\gamma}[p_{S}:p_{S}]$.
Now we go back to the proof. We firstly compute the $\gamma$-cross entropy $d_{\gamma}[p_{S}:p_{\widetilde{S}}]$ for Gaussian probability density functions $p_{S}$ and $p_{\widetilde{S}}$, and then the conclusion directly follows from Szegö’s limit theorem [27, 15]. Since $S$ and $\widetilde{S}$ are positive on $(0,\pi)$, we can show that $\Sigma_{n}(S)$ and $\Sigma_{n}(\widetilde{S})$ are positive definite so that $\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)$ is also positive definite for any $\gamma>0$, which yields $\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)$ is invertible. Then we obtain
$\int_{\mathbb{R}^{n}}p_{S}(\mathbf{x})p_{\widetilde{S}}(\mathbf{x})^{\gamma}\,\mathrm{d}\mathbf{x}$ =$=\left\{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}\right\}^{-\frac{1}{2}}$
$\qquad\quad\int_{\mathbb{R}^{n}}\exp\left(-\frac{1}{2}\mathbf{x}^{\ast}[\Sigma_{n}(S)^{-1}+\gamma\Sigma_{n}(\widetilde{S})^{-1}]\mathbf{x}\right)\,\mathrm{d}\mathbf{x}$
=$=\left\{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}\right\}^{-\frac{1}{2}}$
@@
=$=\left\{\frac{\pi^{n}\det[\Sigma_{n}(\widetilde{S})(\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S))^{-1}\Sigma_{n}(S)]}{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}}\right\}^{\frac{1}{2}}$
=$=\left\{\pi^{-n\gamma}\det[\Sigma_{n}(\widetilde{S})]^{1-\gamma}\det[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]^{-1}\right\}^{\frac{1}{2}}.$
@@
$\int_{\mathbb{R}^{n}}p_{\widetilde{S}}(\mathbf{x})^{1+\gamma}\,\mathrm{d}\mathbf{x}$ =$=\left\{\pi^{-n\gamma}\det[\Sigma_{n}(\widetilde{S})]^{1-\gamma}\det[(1+\gamma)\Sigma_{n}(\widetilde{S})]^{-1}\right\}^{\frac{1}{2}}$
=$=\left\{\pi^{-n\gamma}(1+\gamma)^{-n}\det[\Sigma_{n}(\widetilde{S})]^{-\gamma}\right\}^{\frac{1}{2}}.$
@@
$\qquad+\frac{1}{1+\gamma}\log\left[\pi^{-n\gamma}(1+\gamma)^{-p}\det[\Sigma_{n}(\widetilde{S})]^{-\gamma}\right]$
@@
$\qquad+\left[\frac{1}{\gamma}\log\det[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]-\frac{1}{\gamma(1+\gamma)}\log\det[\Sigma_{n}(\widetilde{S})]\right].$
@@
$2D_{\gamma}[p_{S}:p_{\widetilde{S}}]$ $=\frac{1}{\gamma}\Bigg{[}\log\det[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]-\frac{1}{1+\gamma}\log\det[\Sigma_{n}(\widetilde{S})]$
$\qquad\qquad\qquad\qquad\qquad\qquad-\log(1+\gamma)^{p}-\frac{\gamma}{1+\gamma}\log\det[\Sigma_{n}(S)]\Bigg{]}$
$=\frac{1}{\gamma}\Bigg{[}\log\det\left[\frac{\gamma}{1+\gamma}\Sigma_{n}(S)+\frac{1}{1+\gamma}\Sigma_{n}(\widetilde{S})\right]$
$\qquad\qquad\qquad\qquad-\frac{\gamma}{1+\gamma}\log\det[\Sigma_{n}(S)]-\frac{1}{1+\gamma}\log\det[\Sigma_{n}(\widetilde{S})]\Bigg{]}$
$=\frac{1}{\gamma}\left[\log\det[\Sigma_{n}(\bar{S}^{(\gamma)})]-\frac{\gamma}{1+\gamma}\log\det[\Sigma_{n}(S)]-\frac{1}{1+\gamma}\log\det[\Sigma_{n}(\widetilde{S})]\right],$
where $\bar{S}^{(\gamma)}(\omega):=\frac{\gamma}{1+\gamma}S(\omega)+\frac{1}{1+\gamma}\widetilde{S}(\omega)$. Using Szegö’s limit theorem [27, 15], we conclude
$\lim_{n\to\infty}\frac{2}{n}D_{\gamma}[p_{S}:p_{\widetilde{S}}]$ $=\frac{1}{2\pi\gamma}\int_{-\pi}^{\pi}\left[\log{\bar{S}^{(\gamma)}(\omega)}-\frac{\gamma}{1+\gamma}\log{S(\omega)}-\frac{1}{1+\gamma}\log{\widetilde{S}(\omega)}\right]\,\mathrm{d}\omega$
$=D_{\gamma}[S:\widetilde{S}].$
This completes the proof.

### v4

@@ 4.1 AR(2) model
using the periodogram ($I_{n}$); (4) the minimum Itakura–Saito divergence estimator employing the periodogram ($I_{n}$); and (5) the minimum Itakura–Saito divergence estimator with the smoothed periodogram ($I_{n}^{S}$).
For each divergence minimization problem, the minimization is conducted via the gradient descent algorithm with a fixed learning rate $\gamma=0.005$. We stop the optimization when the optimization step reaches 10000 or when the Euclidean norm of the gradient of each divergence evaluated at each updated estimate is less than $10^{-3}$. For smoothing the periodogram, we apply the modified Daniell smoothers twice, where the length of the first time is $3$ and the length of the second time is $5$.
Tables 1 and 2 summarize the estimation results in the above setup. Tables 1 and 2 suggest that the minimizers of the IS divergence with the periodogram/smoothed periodogram are very sensitive to the trigonometric trends of the observed time series (i.e. the outliers in the frequency domain). On the other hand, the minimizers of the spectral Rényi divergence give stably estimated values and their performance would be substantially better than those of the IS divergence with the periodogram/smoothed periodogram. Thus the estimated values using the spectral Rényi divergence are robust against the outliers in the frequency domain.
Figure 2: Spectral densities with the estimates plugged in for the AR(2) model without any trend. Each panel shows the results for different realizations of the periodogram. The gray curve is the periodogram. The green curve is the smoothed periodogram. The true spectral density is colored in black. The spectral density based on the spectral Rényi divergence is colored in red. The spectral density based on the Itakura–Saito divergence is colored in blue. We set $\theta^{(0),1}$ to the initial value. The initial value $\theta^{(0),\mathrm{best}}$ represents the point that minimizes the objective functions over a grid of the parameter space.
Figure 3: Spectral densities with the estimates plugged in for the AR(2) model with the trend. Each panel shows the results for different realizations of the periodogram. The legend is the same as in Figure 2. The red dot lines represent the points of the outliers. We set $\theta^{(0),1}$ to the initial value.
Figure 4: Histograms of the numbers of iterations to converge in the AR(2) model, where the initial value is $\theta^{(0),2}$. We conducted each simulation 100 times and plotted the histogram. (a) The Itakura–Saito divergence minimization without any trend. (b) The Itakura–Saito divergence minimization with a trend. (c) The $\alpha=0.9$-Rényi divergence minimization without any trend. (d) The $\alpha=0.9$-Rényi divergence minimization with a trend.
Figures 2 and 3 display the estimated spectral densities. In each figure, the solid gray curve represents the periodogram $I_{n}(\omega)$; the solid green curve represents the smoothed periodogram $I_{n}^{S}(\omega)$; the solid black curve represents the spectral density $S^{\mathrm{AR2}}_{\theta}(\omega)$ with the true value $\theta^{*}$ plugged in; the solid blue curve represents the spectral density with the minimum Itakura–Saito divergence estimate plugged in; the solid red curve represents the spectral density with the minimum spectral Rényi divergence ($\alpha=0.9$) estimate plugged in. The dashed curves represent the estimates starting from the initial values that minimize the objective functions, respectively. Here the periodogram is used in these figures. Figures 2 and 3 show that the spectral density $S^{\mathrm{AR2}}_{\widehat{\theta}_{n}}$ with the minimum spectral Rényi divergence ($\alpha=0.9$) (resp. IS divergence) estimate $\widehat{\theta}_{n}$ plugged in fits (resp. does not fit) well to the spectral density $S^{\mathrm{AR2}}_{\theta^{*}}$ with the true value $\theta^{*}$ plugged in, which implies that the minimizers of the IS divergence with the periodogram/smoothed periodogram are very sensitive to the outliers in the frequency domain, but the spectral Rényi divergence is robust against them. These figures also indicate that the Itakura–Saito divergence estimate is highly sensitive to realizations of the periodogram even without any contamination. The grid search for the initial values does not help improve the performance of the Itakura–Saito divergence estimate. In contrast, the spectral Rényi estimate is generally less sensitive to realizations of the periodogram.
Further, Figure 4 displays the histograms of the numbers of iterations required for convergence when the initial value is $\theta^{(0),2}$. For the Rényi-divergence minimization, the vast majority of trials converge within a moderate number of iterations, both with and without outliers. In contrast, for the Itakura–Saito divergence minimization, the convergence behavior deteriorates markedly in the presence of outliers, where essentially all trials reach the maximum number of iterations without convergence. These findings are consistent with the optimization trajectories illustrated in Figure 1.
@@ 4.2 The Brune spectral model with attenuation
Next, consider the Brune spectral model Brune (1970) with attenuation Aki and Richards (1980) for velocity waveforms:
The Brune spectral model $\sigma^{2}/\{1+(\omega/\omega_{c})^{2}\}^{2}$ for displacement waveforms is the spectral density model of Matérn process Matérn (1960) with the spectral decay rate fixed to 2 and often used in the earthquake source estimation Calderoni and Abercrombie (2023); Yoshimitsu et al. (2023). The attenuation factor $\exp(-\omega/Q)$ represents anelasticity of the medium through which the seismic wave propagates. The multiplicative factor $\omega^{2}$ comes from the time derivative that converts a displacement waveform to a velocity waveform.
Figure 5: Spectral densities with the estimates plugged in for the Brune model without any trend. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Figure 6: Spectral densities with the estimates plugged in for the Brune model with the trigonometric trends. The red dashed lines denote the frequencies at which the outliers are injected. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Table 3: Mean values of biases with standard deviations for the Brune model without any trend. Values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
@@
IS with $I_{n}^{S}$ $\theta^{(0),2}$ -0.29 ($\pm$0.30) 1.95 ($\pm$1.89) 1.86 ($\pm$0.98)
R ($\alpha=0.50$) $\theta^{(0),3}$ -0.10 ($\pm$0.04) 0.51 ($\pm$0.13) -0.30 ($\pm$0.06)
R ($\alpha=0.75$) $\theta^{(0),3}$ -0.05 ($\pm$0.06) 0.08 ($\pm$0.19) 0.05 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),3}$ -0.007 ($\pm$0.06) -0.15 ($\pm$0.13) 0.72 ($\pm$0.35)
@@
IS with $I_{n}^{S}$ $\theta^{(0),3}$ -0.29 ($\pm$0.15) 2.09 ($\pm$1.68) 1.50 ($\pm$0.72)
Table 4: Mean values of biases with standard deviations for the Brune model with the trigonometric trends. Values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
@@
with $z_{1}=z_{2}=100$, $A_{1}=\pi/4$ and $A_{2}=\pi/8$. We report the results based on different values of $z_{1}=z_{2}$ in Appendix F. We use three different initial values: $\theta^{(0),1}=(1,0.1,1)$, $\theta^{(0),2}=(1,1,1)$, and $\theta^{(0),3}=(1,2,1)$. Setting $\sigma^{*}$ and $Q^{*}$ as initial values does not impact on the behavior of the minimum spectral Rényi divergence estimates; we report the results based on different initial values $\theta^{(0),4}=(0.1,0.1,0.1)$ and $\theta^{(0),5}=(2,2,2)$ in Appendix F.
Figures 5 and 6 display the estimated spectral densities. In each figure, the following representations are used:
@@
  • the solid blue curve illustrates the spectral density with the minimum Itakura–Saito divergence estimate plugged in;
  • the solid red, salmon pink, and green curves display the spectral densities with the minimum spectral Rényi divergence ($\alpha=0.9,0.75,0.5$) estimates plugged in, respectively.
@@
Tables 3 and 4 summarize the estimation results. Comparing the results based on different initial values, we find that the minimum Itakura–Saito divergence estimates are sensitive to the initial value of the optimization regardless of the existence of the smoother, while the spectral Rényi divergence yields estimation results stable to the choice of the initial value. The minimum Itakura–Saito divergence estimate with the smoothed periodogram performs much worse than that with the periodogram. In the presence of outliers in the frequency domain, the minimum Rényi divergence estimate with $\alpha=0.5$ performs the best regardless of the choice of the initial value in this example.
@@ 5 Discussions
Finally, we discuss three important directions for future research.
The first point is the choice of $\alpha$. The value of $\alpha$ controls the trade-off between the efficiency without the presence of outliers and the robustness even for larger trends. For an objective selection of $\alpha$, constructing an information criterion is one possible direction, but requires sufficient discussions and experiments. In practice, monitoring the behaviors of the minimum spectral Rényi divergence estimates for several values of $\alpha$ (say, $\alpha=0.5,0.75,0.9$) would be a workaround.
The second point is the choice of optimization schemes. In this paper, we employed a simple gradient descent scheme only as a baseline illustrative method, and we do not claim that naive gradient descent is the most appropriate choice in practice. As is well known, the iteration number, tolerance, and step-size selection in naive gradient descent can be highly sensitive and not objectively determined. There are at least two ways to avoid these issues:
  1. 1.
    Use of line-search methods. One can incorporate standard line-search strategies such as Wolfe or Armijo conditions (see Proposition 2 and Remark 3), or alternatively employ recent line-search-free first-order methods Malitsky and Mishchenko (2020); Yagishita and Ito (2025), which often provide more stable convergence behavior without requiring excessively large iteration limits.
  2. 2.
    Use of quasi-Newton methods. Quasi-Newton algorithms such as the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm or L-BFGS Liu and Nocedal (1989) are typically more efficient in terms of optimization and are widely preferred by practitioners. Indeed, in a large-scale application of spectral Rényi divergence minimization to real geophysical data Kano et al. (2025), we adopted the L-BFGS method precisely to avoid the practical difficulties associated with naive first-order methods such as gradient descent.
Constructing a unified theory that accommodates line-search-free first-order methods and quasi-Newton methods would be an important research direction.
The third one is the impact of the choice of an initial value. Numerical experiments presented in Section 4 suggest that the minimum spectral Rényi divergence estimates are robust to the choice of an initial value, while the minimum Itakura–Saito divergence estimate is sensitive to it. These robustness and sensitivity are confirmed in the other examples such as autoregressive models. The structure of the gradient $\mathcal{G}_{\alpha}(\theta\,;\,\widetilde{I}_{n})$ in the comparison with that of $\mathcal{G}_{\mathrm{IS}}(\theta\,;\,\widetilde{I}_{n})$ can give a clue to this, and so detailed analysis of these gradient landscapes would be one of the interesting research directions.
6 Acknowledgements
The authors would like to thank the Associate Editor and the referee for their constructive comments, which improved the quality of the paper. The authors would like to thank Akifumi Okuno, Mirai Tanaka, Kei Kobayashi, Yuta Koike, Tomoyuki Higuchi, Hironori Fujisawa, Junho Yang, Masayuki Kano, and Shotaro Yagishita for their comments. This work is supported by JSPS KAKENHI (19K20222, 21H05205, 21K12067, 23K11024), MEXT (JPJ010217), and “Strategic Research Projects” grant (2022-SRP-13) from ROIS (Research Organization of Information and Systems). The simulation code is available at https://github.com/t-tetsuya/Spectral_Renyi.git.
@@ 6 Acknowledgements
  • Proof of Proposition 3 in Appendix D;
  • Additional comparison of optimization paths in Appendix E; and
  • Additional simulation studies in Appendix F.
@@ Appendix A Proof of Theorem 1
We first introduce the notation used in the proof. For each spectral density function $S$ of a stationary time series $(X_{t})_{t\in\mathbb{Z}}$, we write $\Sigma_{n}(S)$ as the $n\times n$-Toeplitz matrix whose $(s,t)$th element is given by
@@
which corresponds to the covariance matrix of $(X_{1},X_{2},\cdots,X_{n})^{\top}$. Then, the probability density function of $(X_{1},\cdots,X_{n})^{\top}$ is expressed by
@@
Let $\gamma\in(0,\infty)$. First introduce the $\gamma$-cross entropy $d_{\gamma}[\,p_{S}\,:\,p_{\widetilde{S}}\,$] between two probability density functions $p_{S}$ and $p_{\widetilde{S}}$, which are not necessarily Gaussian probability density functions, defined by
@@
Then the $\gamma$-divergence can be written as $G_{\gamma}[\,p_{S}\,:\,p_{\widetilde{S}}\,]=d_{\gamma}[\,p_{S}\,:\,p_{\widetilde{S}}\,]-d_{\gamma}[\,p_{S}\,:\,p_{S}\,]$.
[…]
$\int_{\mathbb{R}^{n}}p_{S}(\mathbf{x})p_{\widetilde{S}}(\mathbf{x})^{\gamma}\,\mathrm{d}\mathbf{x}$ $=\left\{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}\right\}^{-\frac{1}{2}}$
$\qquad\quad\int_{\mathbb{R}^{n}}\exp\left(-\frac{1}{2}\mathbf{x}^{\top}[\Sigma_{n}(S)^{-1}+\gamma\Sigma_{n}(\widetilde{S})^{-1}]\mathbf{x}\right)\,\mathrm{d}\mathbf{x}$
$=\left\{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}\right\}^{-\frac{1}{2}}$
@@
$=\left\{\frac{\pi^{n}\det[\Sigma_{n}(\widetilde{S})(\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S))^{-1}\Sigma_{n}(S)]}{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}}\right\}^{\frac{1}{2}}$
$=\left\{\pi^{-n\gamma}\det[\Sigma_{n}(\widetilde{S})]^{1-\gamma}\det[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]^{-1}\right\}^{\frac{1}{2}}.$
@@
$\int_{\mathbb{R}^{n}}p_{\widetilde{S}}(\mathbf{x})^{1+\gamma}\,\mathrm{d}\mathbf{x}$ $=\left\{\pi^{-n\gamma}\det[\Sigma_{n}(\widetilde{S})]^{1-\gamma}\det[(1+\gamma)\Sigma_{n}(\widetilde{S})]^{-1}\right\}^{\frac{1}{2}}$
$=\left\{\pi^{-n\gamma}(1+\gamma)^{-n}\det[\Sigma_{n}(\widetilde{S})]^{-\gamma}\right\}^{\frac{1}{2}}.$
@@
$\qquad+\frac{1}{1+\gamma}\log\left[\pi^{-n\gamma}(1+\gamma)^{-n}\det[\Sigma_{n}(\widetilde{S})]^{-\gamma}\right]$
@@
[…]
@@
[…]
$\qquad\qquad\qquad\qquad\qquad\qquad-\log(1+\gamma)^{n}-\frac{\gamma}{1+\gamma}\log\det[\Sigma_{n}(S)]\Bigg]$
[…]
where $\bar{S}^{(\gamma)}(\omega):=\frac{\gamma}{1+\gamma}S(\omega)+\frac{1}{1+\gamma}\widetilde{S}(\omega)$. Using Szegö’s limit theorem Szegö (1920); Kolmogorov (1941), we conclude
$\lim_{n\to\infty}\frac{2}{n}G_{\gamma}[\,p_{S}\,:\,p_{\widetilde{S}}\,$] $=\frac{1}{2\pi\gamma}\int_{-\pi}^{\pi}\left[\log{\bar{S}^{(\gamma)}(\omega)}-\frac{\gamma}{1+\gamma}\log{S(\omega)}-\frac{1}{1+\gamma}\log{\widetilde{S}(\omega)}\right]\,\mathrm{d}\omega$
$=D_{\alpha}[\,S\,:\,\widetilde{S}\,$],
where $\alpha:=(1+\gamma)^{-1}$. This completes the proof.